OpenStudy (anonymous):

A coin is tossed 100 times resulting in 46 heads and 54 tails. The same coin is tossed 1000 times resulting in 495 heads and 505 tails. What is the theoretical probability of getting heads with this coin? A. 49.2% B. 49.5% C. 50% D. 46%

OpenStudy (kropot72):

Out of 1100 tosses heads came up 45 + 495 = 541 times and tails came up 54 + 505 = 559 times. Therefore the experimental probability of getting heads is 541/1100. However if the coin is fair, the theoretical probability of getting heads is 1/2 = 0.5.

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)

OpenStudy (kropot72):

@dbrown13 Do you understand which answer option to choose now?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Which one did you choose?

OpenStudy (anonymous):

C.

OpenStudy (kropot72):

Your choice is correct :)